# A chaotic lattice field theory in two dimensions

### v2

@@ IX.2 Periodic state stability
where $u^{(\alpha)}(\mathsf{k})$ are band-index $\alpha=1,2,\cdots,{N}_{\mathbb{A}}$ labelled distinct $\mathcal{L}_{\mathbb{A}}$-periodic functions, and the continuous wave numbers $\mathsf{k}$ are restricted to a Brillouin zone $\mathbb{B}$. In solid-state physics, eigenstates Eq. (110) are known as Bloch states. [124] In mechanics they are called Floquet modes. [145]
(a)   Refer to caption
(b)   Refer to caption
Figure 11: (Color online)    Square spatiotemporal lattice orbit Jacobian operator spectra, as functions of the wave vectors $(k_{1},k_{2})$. For time and space-reflection and interchange invariant periodic states the spectra are $k_{1}\to-k_{1}$, $k_{2}\to-k_{2}$ and $k_{1}\leftrightarrow k_{2}$ symmetric. (a) The steady state $\Lambda(\mathsf{k})$ Bloch band Eq. (106) as a function of the wave vector $\mathsf{k}$, plotted for $\mu^{2}=1$ value. Black dots are eigenvalues of the orbit Jacobian matrix of periodic states over primitive cell with periodicity $8\!\times\!8]_{0}$. (b) The two-dimensional $\phi^{4}$ lattice field theory spectra of the Bravais lattice $\mathcal{L}_{[2\!\times\!1]_{0}}$ periodic state Eq. (198), plotted for $\mu^{2}=5$ value. Black dots are eigenvalues of the orbit Jacobian matrix of a $6\!\times\!4]_{0}$ primitive cell tiled by 12 repeats of a prime $2\!\times\!1]_{0}$ periodic state, with $\Lambda_{\pm}(\mathsf{k})$ Bloch bands computed in Appendix C.2.
@@
where $\Lambda_{p,\alpha}(\mathbf{k})$ is the eigenvalue of the prime orbit $p$ orbit Jacobian operator on the $\alpha$-th eigenvalue band, corresponding to the eigenstate $\varphi^{(\alpha)}(\mathsf{k})$ in Eq. (110). This prime orbit stability exponent formula is, in the spirit of Sec. VI.2, the generalization of the steady state stability exponent Eq. (106) to all periodic states.
@@
(Continued from Sec. VIII.3.) Consider the simplest not-steady solution of $\phi^{3}$ theory Eq. (45), its period-2 periodic state. The orbit Jacobian operator Eq. (69) of the period-2 prime orbit Eq. (196) is invariant under translations of period L=2{L}=2, so its first Brillouin zone Eq. (104) is $(-\pi/2,\pi/2]$. This orbit Jacobian operator has two bands (for details, see Appendix C.1):
@@
Have a look back at the period-6 primitive cell of Sec. VIII.3, with the prime periodic state $\mathsf{\Phi}_{01}$ repeated three times. The wave numbers $\mathsf{k}$ are constrained to the period-6 reciprocal lattice of, taking values $\mathsf{k}=-\pi/3,0,\pi/3$ within the first Brillouin zone. They are embedded into the two continuous eigenvalue bands Eq. (115), corresponding to the Eq. (101) eigenvalues $\Lambda_{-1,\pm}$, $\Lambda_{0,\pm}$ and $\Lambda_{1,\pm}$ respectively, as illustrated in Fig. 9 (b).
@@
(Continued from Sec. VIII.3.) Figure 11 (b) shows the two eigenvalue bands of a two-dimensional $\phi^{4}$ $2\!\times\!1]_{0}$ periodic state, plotted over the two-dimensional first Brillouin zone Eq. (104) $k_{1}\in(-\pi/2,\pi/2]$, $k_{2}\in(-\pi,\pi]$. For details, see Appendix C.2.
In summary, while a prime periodic state and its repeats have different orbit Jacobian matrices spectra, they share the same orbit Jacobian operator stability exponent, determined by integration over continuous Bloch bands. Inspection of Figs. 9 and 11 makes it clear what these different spectra are: orbit Jacobian matrix spectra are discrete approximations to orbit Jacobian operator continuous Bloch bands, with ‘cords’ approximation errors decreasing as the primitive cell volume increases. (For the convergence rate of such approximations, see shadowing Sec. XI). The wonderful property of the stability exponent ${\lambda}_{p}$ computed over the infinite spacetime (as opposed to the stability exponent computed over a finite primitive cell) is that it is additive for prime periodic states repeated over larger primitive cells.
X Periodic orbit theory
Now that we know how to enumerate all Bravais lattices $\mathcal{L}_{\mathbb{A}}$ (Sec. V), determine all periodic states over each (Sec. VI), and compute the weight of each periodic state (Sec. IX), we can combine all of that into one generating function sum of a very simple form, for any deterministic field theory, in any spacetime dimension:
@@ X Periodic orbit theory
Notation $t_{c}$’ is a vestige of referring to this weight in the time-evolution periodic orbit theory as the ‘local trace’ (see ChaosBook  sect. 18.2). [146] Indeed, much of the time-evolution periodic orbit theory developed in ChaosBook generalizes to the multi-periodic, spatiotemporal deterministic field theory, with time period $T_{c}$ replaced by the spatiotemporal volume ${N}_{c}$. Square brackets $\cdots$] in quantities such as $Z[\beta,z]$ are a here to remind us that we are dealing with spatiotemporal field theories, [147] not with a few degrees-of-freedom evolving forward in time (Sec. II).
X.1 Chaotic field theory
Ergodic theory of time-evolving dynamical systems is a rich subject. In this series of papers we stay within its most robust corner that we refer to as the ‘chaotic field theory’. We say that a deterministic field theory is chaotic if (1) all of its periodic states are unstable, i.e., the stability exponent, Eq. (111), is strictly positive, ${\lambda}_{c}>0$, for every deterministic solution $\mathsf{\Phi}_{c}$, and (2) the number of periodic states $|c|_{\mathbb{A}}$ grows exponentially with the primitive cell volume ${N}_{\mathbb{A}}$, with (3) periodic states are ‘shadowed’ by lower volume periodic states (Sec. XI).
To understand where the ‘generating function variable’ $z$ in Eq. (116) comes from, consider $Z_{\mathbb{A}}[\beta]$, the primitive cell $\mathbb{A}$ partition sum Eq. (33) over all periodic states $\mathsf{\Phi}_{c}$ of periodicity $\mathcal{L}_{\mathbb{A}}$. Their number $|c|_{\mathbb{A}}$ is the number of admissible mosaics (Sec. II.3), with the mean of the log of the number of periodic states per lattice site given by $h_{\mathbb{A}}=\frac{1}{{N}_{\mathbb{A}}}\ln|c|_{\mathbb{A}}\,.$
@@ X.1 Chaotic field theory
The exact value of $h_{\mathbb{A}}$ might require a calculation, and evaluation of the expectation value of system’s entropy $h$ will require the full machinery of the periodic orbit theory developed here in Sec. X.6, but all we need to ensure that the theory is spatiotemporally chaotic is that all $h_{\mathbb{A}}$ are strictly positive.
@@
as illustrated by Figs. 9 and 11 and the anti-integrable limit calculations Eq. (67) and Eq. (108), so:
@@
a sum over all ‘geometries’. Exponential bound Eq. (120) ensures that the sum is convergent for sufficiently small generating function variable $z$. With the stability exponent evaluated over the reciprocal lattice, as in Sec. IX, this is just the deterministic partition sum Eq. (116) arranged as a series in $z^{{N}}$. Our first task will be to determine the largest $z$ for which the deterministic partition sum is convergent, Sec. X.5.
In this paper we utilize only the translation group $T$ symmetries, Eq. (70) (see Sec. V), and focus on the case of a two-dimensional square lattice. Translations stratify the deterministic field theory, Eq. (116), into prime orbits[149, 150, 20] periodic states that are not repeats of a shorter period state (Secs. I.3 and VI.1). To assemble the deterministic partition sum over all periodic states, we reorganize it by first determining all prime orbits $\mathsf{\Phi}_{p}$, and summing over their repeats (Sec. VI.2). Then the deterministic partition sum Eq. (116) takes form
@@
X.2 Steady state partition sum
A steady state $\phi_{z}={\phi}$ is a prime periodic state whose primitive cell $1\!\times\!1]_{0}$} is the unit hypercube Eq. (5) (Sec. VI.2). A periodic state is then obtained by tiling any larger primitive cell by repeats of steady state $\phi$}, one such periodic state for every Bravais sublattice.
For a one-dimensional, temporal Bravais lattice, the deterministic partition sum Eq. (116) is very simple. There is a non-prime periodic state for every repeat primitive cell $\mathbb{A}$ of period $r$, all of them with orbits of period 1 (Sec. I.3), so thanks to repeat weights multiplicativity Eq. (113), the contribution of $r$th repeat to the partition sum Eq. (116) is $t_{p}^{r}\,=\,\left({e}^{\beta\cdot{a}_{p}-{\lambda}_{p}}{z}\right)^{r},$ where r{r} is the volume of the period r{r} primitive cell. So, in one dimension the steady state partition sum is a geometric series,
@@ X.2 Steady state partition sum
So for a two-dimensional, spatiotemporal steady state (see Sec. VI) there is a non-prime periodic state for every two-dimensional repeat primitive cell $r_{1}\!\times\!r_{2}]_{s}$} constructed in Sec. V.1, with $r_{1}$ copies of lattice site field $\phi{}$ horizontally, $r_{2}$ copies vertically, with weights, as stated in Eq. (112), independent of the tilt $0\leq s<r_{1}$ Eq. (79),
@@
was first studied by Euler, with $\sigma({n})$ known as the Euler sum-of-divisors function.
X.3 Prime orbit partition sum
As explained in Sec. VI.2, every prime periodic state is morally a steady state, except that now the number of periodic states in the prime orbit is its primitive cell volume ${N}_{p}$. Hence in one dimension the prime orbit $p$ deterministic partition sum, including steady states Eq. (123) as the ${N}_{p}=1$ cases, is simply
@@ X.3 Prime orbit partition sum
The formula has form of the ‘deterministic trace formula’ (see ChaosBook  Eq. (21.24)), [151] with a crucial difference: here the weight $t_{p}$ is the exact, infinite Bravais lattice weight, while the Ruelle dynamical zeta function weight $t_{p}$ is an approximate weight.
@@
X.4 Spatiotemporal zeta functions
We have a deterministic partition sum Eq. (116) of stunning simplicity. Still, we can do better, by taking into account symmetries (Sec. V) of the theory.
@@ X.4 Spatiotemporal zeta functions
Why? (1) The deterministic partition sum Eq. (116) is a redundant sum over all periodic states, redundant as their weights depend only on prime orbits, which a zeta function counts only once per orbit. (2) Every periodic state weight contributes to the deterministic partition sum with a positive weight. Zeta functions are smarter, they exploit the key property of ergodic trajectories that they are shadowed by shorter trajectories (Sec. XI), with convergence of periodic states averaging formulas improved by shadowing cancellations. (3) Zeta functions have better analyticity properties, with divergence of deterministic partition sum Eq. (121) corresponding to the leading zero of deterministic zeta function.
In one spatiotemporal dimension, the deterministic zeta function is a product over all prime orbits, of form margin: [Uncaptioned image] margin: [Uncaptioned image]
@@
X.5 Evaluation of zeta functions
@@
In dynamical systems theory, the rate at which trajectories leave an open system per unit time is called escape rate (see Ref. 154 and ChaosBook  Eq. (1.3)). We put ‘reject rate’ into quotations here, as in spatiotemporal theory there is no escape in time - the exponent is a characterization of the non–wandering set, the state space set formed by the deterministic solutions. We evaluate it for the temporal cat in Appendix B.5, and for deterministic $\phi^{3}$ and $\phi^{4}$ theories in companion paper III. [17]
Much is known about the two-spatiotemporal dimensions zeta function, Eq. (129), as for each prime orbit $1/\zeta_{p}$ is the Euler function $\phi(t_{p})$,
@@
X.6 Periodic states averaging formula
While each primitive cell probability density Eq. (25) is easily correctly normalized, there is no natural ‘overall’ probability normalization for the generating function sum over all periodic states, the deterministic partition sum Eq. (116). For one-dimensional, temporally evolving systems, ChaosBook  sect. 23.5 solves this problem by the method of ‘cycle averaging’. [153] We now generalize this method to the higher-dimensional, spatiotemporal field theories.
@@ X.6 Periodic states averaging formula
is now given by the periodic states averaging formula margin: [Uncaptioned image] margin: [Uncaptioned image]
@@
While power series expansions in $z$ of functions such as the Euler function, Eq. (135), do not converge very well, the theory of doubly-periodic elliptic functions suggests other, more powerful methods to evaluate such functions. The Euler function can be expressed as the Dedekind eta function $\eta(\tau)$,
@@
The derivatives required for the evaluation of expectation values, Eq. (138), have their own elliptic functions representations. For example, the logarithmic derivative of Dedekind eta is known as the Weierstrass zeta function,
@@
Next –the beauty of the periodic orbit theory of chaos– due to the shadowing of longer periods unstable periodic states by shorter periods ones, the smallest periodicites periodic states dominate, the longer ones come in only as corrections. The convergence periodic states-expansions is accelerated by shadowing of long orbits by shorter periodic orbits. [49] Are $d$-dimensional tori (primitive cells) periodic states also shadowed by smaller tori periodic states? In Sec. XI we check numerically that spatiotemporal cat periodic states that share finite spatiotemporal mosaics indeed shadow each other to exponential precision.
XI Shadowing

### v3

@@ X.2 Periodic state stability
where $u^{(\alpha)}(\mathsf{k})$ are band-index $\alpha=1,2,\cdots,{N}_{\mathbb{A}}$ labeled distinct $\mathcal{L}_{\mathbb{A}}$-periodic functions, and the continuous wave numbers $\mathsf{k}$ are restricted to a Brillouin zone $\mathbb{B}$. In solid-state physics, eigenstates Eq. (110) are known as Bloch states. [130] In mechanics they are called Floquet modes. [151]
@@
where $\Lambda_{p,\alpha}(\mathbf{k})$ is the eigenvalue of the prime orbit $p$ orbit Jacobian operator on the $\alpha$-th eigenvalue band, corresponding to the eigenstate $\varphi^{(\alpha)}(\mathsf{k})$ in Eq. (110). This prime orbit stability exponent formula is, in the spirit of Sec. VII.2, the generalization of the steady state stability exponent Eq. (106) to all periodic states.
@@
(This continues from Sec. IX.3.) Consider the simplest not-steady solution of $\phi^{3}$ theory Eq. (45), its period-2 periodic state. The orbit Jacobian operator Eq. (69) of the period-2 prime orbit Eq. (196) is invariant under translations of period L=2{L}=2, so its first Brillouin zone Eq. (104) is $(-\pi/2,\pi/2]$. This orbit Jacobian operator has two bands (for details, see Appendix C.1):
@@
Have a look back at the period-6 primitive cell of Sec. IX.3, with the prime periodic state $\mathsf{\Phi}_{01}$ repeated three times. The wave numbers $\mathsf{k}$ are constrained to the period-6 reciprocal lattice of, taking values $\mathsf{k}=-\pi/3,0,\pi/3$ within the first Brillouin zone. They are embedded into the two continuous eigenvalue bands Eq. (115), corresponding to the Eq. (101) eigenvalues $\Lambda_{-1,\pm}$, $\Lambda_{0,\pm}$ and $\Lambda_{1,\pm}$ respectively, as illustrated in Fig. 11 (b).
@@
(This continues from Sec. IX.3.) Figure 10 (b) shows the two eigenvalue bands of a two-dimensional $\phi^{4}$ $2\!\times\!1]_{0}$ periodic state, plotted over the two-dimensional first Brillouin zone Eq. (104) $k_{1}\in(-\pi/2,\pi/2]$, $k_{2}\in(-\pi,\pi]$. For details, see Appendix C.2.
In summary, while a prime periodic state and its repeats have different orbit Jacobian matrices spectra, they share the same orbit Jacobian operator stability exponent, determined by integration over continuous Bloch bands. Inspection of Figs. 11 and 10 makes it clear what these different spectra are: orbit Jacobian matrix spectra are discrete approximations to orbit Jacobian operator continuous Bloch bands, with ‘cords’ approximation errors decreasing as the primitive cell volume increases. (For the convergence rate of such approximations, see shadowing Sec. XII). The wonderful property of the stability exponent ${\lambda}_{p}$ computed over the infinite spacetime (as opposed to the stability exponent computed over a finite primitive cell) is that it is additive for prime periodic states repeated over larger primitive cells.
XI Periodic orbit theory
Now that we know how to enumerate all Bravais lattices $\mathcal{L}_{\mathbb{A}}$ (Sec. VI), determine all periodic states over each (Sec. VII), and compute the weight of each periodic state (Sec. X), we can combine all of that into one generating function sum of a very simple form, for any deterministic field theory, in any spacetime dimension:
@@ XI Periodic orbit theory
Notation $t_{c}$’ is a vestige of referring to this weight in the time-evolution periodic orbit theory as the ‘local trace’ (see ChaosBook  sect. 18.2). [152] Indeed, much of the time-evolution periodic orbit theory developed in ChaosBook generalizes to the multi-periodic, spatiotemporal deterministic field theory, with time period $T_{c}$ replaced by the spatiotemporal volume ${N}_{c}$. Square brackets $\cdots$] in quantities such as $Z[\beta,z]$ are a here to remind us that we are dealing with spatiotemporal field theories, [153] not with a few degrees-of-freedom evolving forward in time (Sec. III).
XI.1 Chaotic field theory
Ergodic theory of time-evolving dynamical systems is a rich subject. In this series of papers we stay within its most robust corner that we refer to as the ‘chaotic field theory’. We say that a deterministic field theory is chaotic if (1) all of its periodic states are unstable, i.e., the stability exponent, Eq. (111), is strictly positive, ${\lambda}_{c}>0$, for every deterministic solution $\mathsf{\Phi}_{c}$, and (2) the number of periodic states $|c|_{\mathbb{A}}$ grows exponentially with the primitive cell volume ${N}_{\mathbb{A}}$, with (3) periodic states are ‘shadowed’ by lower volume periodic states (Sec. XII).
To understand where the ‘generating function variable’ $z$ in Eq. (116) comes from, consider $Z_{\mathbb{A}}[\beta]$, the primitive cell $\mathbb{A}$ partition sum Eq. (33) over all periodic states $\mathsf{\Phi}_{c}$ of periodicity $\mathcal{L}_{\mathbb{A}}$. Their number $|c|_{\mathbb{A}}$ is the number of admissible mosaics (Sec. III.3), with the mean of the log of the number of periodic states per lattice site given by $h_{\mathbb{A}}=\frac{1}{{N}_{\mathbb{A}}}\ln|c|_{\mathbb{A}}\,.$
@@ XI.1 Chaotic field theory
The exact value of $h_{\mathbb{A}}$ might require a calculation, and evaluation of the expectation value of system’s entropy $h$ will require the full machinery of the periodic orbit theory developed here in Sec. XI.6, but all we need to ensure that the theory is spatiotemporally chaotic is that all $h_{\mathbb{A}}$ are strictly positive.
@@
as illustrated by Figs. 11 and 10 and the anti-integrable limit calculations Eq. (67) and Eq. (108), so:
@@
a sum over all ‘geometries’. Exponential bound Eq. (120) ensures that the sum is convergent for sufficiently small generating function variable $z$. With the stability exponent evaluated over the reciprocal lattice, as in Sec. X, this is just the deterministic partition sum Eq. (116) arranged as a series in $z^{{N}}$. Our first task will be to determine the largest $z$ for which the deterministic partition sum is convergent, Sec. XI.5.
In this paper we utilize only the translation group $T$ symmetries, Eq. (70) (see Sec. VI), and focus on the case of a two-dimensional square lattice. Translations stratify the deterministic field theory, Eq. (116), into prime orbits[155, 156, 20] periodic states that are not repeats of a shorter period state (sections II.3 and VII.1). To assemble the deterministic partition sum over all periodic states, we reorganize it by first determining all prime orbits $\mathsf{\Phi}_{p}$, and summing over their repeats (Sec. VII.2). Then the deterministic partition sum Eq. (116) takes form
@@
XI.2 Steady state partition sum
A steady state $\phi_{z}={\phi}$ is a prime periodic state whose primitive cell $1\!\times\!1]_{0}$} is the unit hypercube Eq. (5) (Sec. VII.2). A periodic state is then obtained by tiling any larger primitive cell by repeats of steady state $\phi$}, one such periodic state for every Bravais sublattice.
For a one-dimensional, temporal Bravais lattice, the deterministic partition sum Eq. (116) is very simple. There is a non-prime periodic state for every repeat primitive cell $\mathbb{A}$ of period $r$, all of them with orbits of period 1 (Sec. II.3), so thanks to repeat weights multiplicativity Eq. (113), the contribution of $r$th repeat to the partition sum Eq. (116) is $t_{p}^{r}\,=\,\left({e}^{\beta\cdot{a}_{p}-{\lambda}_{p}}{z}\right)^{r},$ where r{r} is the volume of the period r{r} primitive cell. So, in one dimension the steady state partition sum is a geometric series,
@@ XI.2 Steady state partition sum
So for a two-dimensional, spatiotemporal steady state (see Sec. VII) there is a non-prime periodic state for every two-dimensional repeat primitive cell $r_{1}\!\times\!r_{2}]_{s}$} constructed in Sec. VI.1, with $r_{1}$ copies of lattice site field $\phi{}$ horizontally, $r_{2}$ copies vertically, with weights, as stated in Eq. (112), independent of the tilt $0\leq s<r_{1}$ Eq. (79),
@@
was first studied by Euler, with $\sigma({n})$ known as the Euler sum-of-divisors function. [157]
XI.3 Prime orbit partition sum
As explained in Sec. VII.2, every prime periodic state is morally a steady state, except that now the number of periodic states in the prime orbit is its primitive cell volume ${N}_{p}$. Hence in one dimension the prime orbit $p$ deterministic partition sum, including steady states Eq. (123) as the ${N}_{p}=1$ cases, is simply
@@ XI.3 Prime orbit partition sum
The formula has form of the ‘deterministic trace formula’ (see Eq. (21.24) in Ref. 158) with a crucial difference: here the weight $t_{p}$ is the exact, infinite Bravais lattice weight, while the Ruelle dynamical zeta function weight $t_{p}$ is an approximate weight.
@@
XI.4 Spatiotemporal zeta functions
We have a deterministic partition sum Eq. (116) of stunning simplicity. Still, we can do better, by taking into account symmetries (Sec. VI) of the theory.
@@ XI.4 Spatiotemporal zeta functions
Why? (1) The deterministic partition sum Eq. (116) is a redundant sum over all periodic states, redundant as their weights depend only on prime orbits, which a zeta function counts only once per orbit. (2) Every periodic state weight contributes to the deterministic partition sum with a positive weight. Zeta functions are smarter, they exploit the key property of ergodic trajectories that they are shadowed by shorter trajectories (Sec. XII), with convergence of periodic states averaging formulas improved by shadowing cancellations. (3) Zeta functions have better analyticity properties, with divergence of deterministic partition sum Eq. (121) corresponding to the leading zero of deterministic zeta function.
In one spatiotemporal dimension, the deterministic zeta function is a product over all prime orbits, of form [160] margin: [Uncaptioned image] margin: [Uncaptioned image]
@@
XI.5 Evaluation of zeta functions
@@
In dynamical systems theory, the rate at which trajectories leave an open system per unit time is called escape rate (see Ref. 162 and ChaosBook  Eq. (1.3)). We put ‘reject rate’ into quotations here, as in spatiotemporal theory there is no escape in time - the exponent is a characterization of the non–wandering set, the state space set formed by the deterministic solutions. We evaluate it for the temporal cat in Appendix B.5, and for deterministic $\phi^{3}$ and $\phi^{4}$ theories in companion paper III. [17]
Much is known about the two-spatiotemporal dimensions zeta function, Eq. (129), as for each prime orbit $1/\zeta_{p}$ is the Euler function $\phi(t_{p})$[163]
@@
XI.6 Periodic states averaging formula
While each primitive cell probability density Eq. (25) is easily correctly normalized, there is no natural ‘overall’ probability normalization for the generating function sum over all periodic states, the deterministic partition sum Eq. (116). For one-dimensional, temporally evolving systems, Sec. 23.5 of Ref. 161 solves this problem by the method of ‘cycle averaging’. We now generalize this method to the higher-dimensional, spatiotemporal field theories.
@@ XI.6 Periodic states averaging formula
is now given by the periodic states averaging formula [160] margin: [Uncaptioned image] margin: [Uncaptioned image]
@@
While power series expansions in $z$ of functions such as the Euler function, Eq. (135), do not converge very well, the theory of doubly-periodic elliptic functions suggests other, more powerful methods to evaluate such functions. The Euler function can be expressed as the Dedekind eta function $\eta(\tau)$[166]
@@
The derivatives required for the evaluation of expectation values, Eq. (138), have their own elliptic functions representations. For example, the logarithmic derivative of Dedekind eta is known as the Weierstrass zeta function, [167]
@@
Next –the beauty of the periodic orbit theory of chaos– due to the shadowing of longer periods unstable periodic states by shorter periods ones, the smallest periodicites periodic states dominate, the longer ones come in only as corrections. The convergence periodic states-expansions is accelerated by shadowing of long orbits by shorter periodic orbits. [49] Are $d$-dimensional tori (primitive cells) periodic states also shadowed by smaller tori periodic states? In Sec. XII we check numerically that spatiotemporal cat periodic states that share finite spatiotemporal mosaics indeed shadow each other to exponential precision.
XII Shadowing